• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 09:38:55 on August 2,2005

    Entry number 150425

    keyword=cavity monitor update

    Update on the new cavity monitors from a recent production run 4034.

    They have recovered from their flatlining (halog 150194).
    Fig1 shows the cavity BCM vs bcm1. cavity 2 & 3 are single
    valued, but cavity 1 has some "parallel drops" away from the line (?!)
    Fig 2 shows a illustrative set of correlations of cavities 2 & 3
    (those near target) to the striplines. First comment is that
    even comparing 4AX to 4BX (stripline to stripline) is not trivial
    since there are presumably strong beta functions. Considering also
    the previously observed X-Y "coupling" (halog 150310), the
    "cloudiness" of the correlation is perhaps NOT a bad thing.
    However I am still bothered by the nonlinearity in halog 150310.
    There is no evidence of strong linearity in this recent run I
    checked, however (though the range of motion was small).

    To summarize where we are:

    1. We're not finished with the gain adjustments as we
    are not making good use of the ADC dynamic range.

    2. I don't think Musson finished the phase adjustments for X (some
    problems last Friday).

    3. The cavity signals "died" on the weekend and then recovered
    mysteriously. We need to know why and what to do about it in
    the future.

    4. There is evidence of nonlinearity in cavity 2X, see halog 150310.

    5. Understand better the correlation to striplines.


    FIGURE 1

    FIGURE 2